OpenStudy (anonymous):

what is the lcd of ... 4y/5 -y/4+7/10=7/2

OpenStudy (anonymous):

Do you know what your denominators are?

OpenStudy (anonymous):

y=4.845

OpenStudy (anonymous):

can you just answer the question instead of being a douch?

OpenStudy (anonymous):

@emekaojei he's not asking for a solution, he's asking about the LCD.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

This forum isn't put here to give out answers, douche.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@emekaojei, I wasn't calling you a douche. I was responding in kind to the other mushbrain who decided I must be a douche because I won't give him the answer.

OpenStudy (anonymous):

ok sir

OpenStudy (anonymous):

:-)

OpenStudy (whpalmer4):

\[\frac{4y}{5} -\frac{y}{4}+\frac{7}{10}=\frac{7}{2}\] You need to find the LCD (lowest common denominator) of 2,4,5,10. This is the same as finding the lowest common multiple. Straightforward way of doing that is to list the multiples of each one until you find the first number that appears in each list: 2: 2 4 6 8 10 12 14 16 18 20 22 24 26 28 30 32 4: 4 8 12 16 20 24 28 32 5: 5 10 15 20 25 30 10: 10 20 30 Or, you can apply some intelligence. Factor each number into its prime factors: 2: 2 4: 2*2 5: 5 10: 2*5 Now collect the highest power of each unique factor, and multiply them together. We've got 2's, and we've got 5's here. The highest power of 2 is 2^2=4, and the highest power of 5 is 5. 4*5 = 20. If you look back at the list, you'll see that 20 is the first number to appear in each of the lists.

OpenStudy (whpalmer4):

Actually, I suppose you don't have to include 2 in the LCD computation for this problem, but as it is a factor of one of the other denominators, it doesn't affect the result.